# Generative AI-enabled Wireless Communications for Robust Low-Altitude Economy Networking

### v1

@@ Abstract
Low-Altitude Economy Networks (LAENets) have emerged as significant enablers of social activities, offering low-altitude services such as the transportation of packages, groceries, and medical supplies. Unlike traditional terrestrial networks, LAENets are characterized by control mechanisms and ever-changing operational factors, which make them more complex and susceptible to vulnerabilities. As applications of LAENet continue to expand, robustness of these systems becomes crucial. In this paper, we investigate a novel application of Generative Artificial Intelligence (GenAI) to improve the robustness of LAENets. We conduct a systematic analysis of robustness requirements for LAENets, complemented by a comprehensive review of robust Quality of Service (QoS) metrics from the wireless physical layer perspective. We then investigate existing GenAI-enabled approaches for robustness enhancement. This leads to our proposal of a novel diffusion-based optimization framework with a Mixture of Expert (MoE)-transformer actor network. In the robust beamforming case study, the proposed framework demonstrates its effectiveness by optimizing beamforming under uncertainties, achieving a more than 44% increase in the worst-case achievable secrecy rate. These findings highlight the significant potential of GenAI in strengthening LAENet robustness.
@@ I Introduction
Recently, advancements in wireless communications and manufacturing technologies have propelled the development of drones and other aerial devices, greatly expanding their applications in low-altitude airspace. Enabled by these technologies, the Low-Altitude Economy (LAE) has emerged as a significant sector focusing on social opportunities within low-altitude airspace (defined as elevations up to 1,000 meters). The LAE aims to harness low-flying equipment for various social activities, including air traffic management, transportation, and surveillance, to organize the Low-Altitude Economy Networks (LAENet). These networks are essential as they provide critical services such as communication, computation, storage, and sensing [1].
@@
However, the unique low-altitude environment and the absence of unified management regulations have introduced several uncertainties to the design of LAENet. These uncertainties, including variable weather conditions, air traffic, user demand, and emergency situations, significantly heighten the requirement for LAENet devices to complete tasks in dynamic environments [1]. Therefore, robustness has been recognized as a Key Performance Indicator (KPI) that goes beyond efficiency (speed) and effectiveness (accuracy) for LAENet. Robustness refers to the ability to maintain the qualities of service (QoS) of LAENets despite the influence of uncertainties [2]. This recognition guides the LAENet’s evolution, enabling innovation and commercial activity. For example, the tech giant Uber’s drone delivery utilizes advanced artificial intelligence algorithms, GPS systems, and sensors to avoid obstacles under different weather conditions, ensuring maximum delivery safety. However, due to the complexity of the airspace environment, the high degrees of freedom, and the strong dynamics of aerial devices, building a robust wireless network platform for LAE applications is challenging.
@@
Enhancing the robustness of the wireless physical layer requires addressing uncertainty as a critical issue. Zero-forcing techniques are applicable for accurate signal detection [3]. Deep learning-based joint source-channel coding can provide a reliable method for data transmission [4]. However, both numerical algorithms and traditional learning methods face several challenges, including heavy dependence on training data, and susceptibility to unknown environments, which hinder their effectiveness in handling uncertainty. Generative artificial intelligence (GenAI) is an emerging technology that works to analyze data distribution. GenAI has demonstrated strong capabilities in enhancing wireless communication, including physical and network layers, through cross-dimensional data processing and generation [5]. Based on the achievements in wireless communication, there are several potential advantages of using GenAI to eliminate uncertainty and construct robust LAENets. GenAI can generate additional signal data, which is difficult to obtain in extreme scenarios, to handle diverse interference patterns. Additionally, GenAI can generate clean data, thereby removing the noise uncertainty in signals. Moreover, GenAI can learn patterns from channel transmission data to realistically simulate fading effects and provide precise uncertainty estimates.
@@
Motivated by these analyses, this paper aims to investigate the incorporation of GenAI into the development of robust LAENets. The main contributions are summarized as follows:
  • We provide an overview of LAENets and their applications, offering a detailed analysis of the robustness within LAENet. Furthermore, we introduce various GenAI models and their applications in wireless communications toward LAENets.
  • From the perspective of the physical layer, we explore how GenAI enhances the robustness of wireless communication. Moreover, we present several robustness metrics and discuss the benefits of implementing GenAI models to support the construction of LAENets.
@@
II The overall of Robust LAENet and GenAI
@@
Figure 1: An overall of GenAI-enabled robust LAENet. Part A illustrates the architecture of LAENet from transportation, communication, and sensing. Part B introduces three aspects of robustness in LAENets. Part C presents a comparison between GenAI and deterministic AI. Part D demonstrates important GAI-enabled applications, including channel modeling, channel estimation, and beam alignment.
@@ II-A Conceptions and Applications of LAENet
The LAE relies on the cooperation of edge equipment and low-altitude communication networks to facilitate activities, which covers three primary aspects (shown in Fig. 1 Part A):
@@ II-B The Significance of Robust LAENet
  • Safety: Safety in LAENets involves ensuring the protection of equipment, operators, and service recipients during operations with low-altitude devices. Ensuring QoS in these scenarios means that safety measures must consistently perform effectively, even when faced with challenges such as adverse weather, signal interference, or unexpected obstacles.
  • Security and Confidentiality: Security and confidentiality concern LAENet’s ability to maintain secure and reliable low-altitude communications while ensuring QoS amidst various uncertainties. This involves resisting external attacks, protecting against data breaches, and preventing unauthorized access, all without compromising the performance required for different applications.
@@ II-C Robustness Considerations for LAENet
To effectively evaluate and enhance the robustness of LAENet, it is necessary to introduce robust QoS metrics in various applications. As shown in Fig. 2 Part A, the key applications and corresponding indicators are as follows:
@@
Besides robust indicators, the design and deployment of robust LAENets must account for these uncertainties to meet operational requirements. We present several optimization frameworks that offer increasing levels of robustness, including deterministic, stochastic, chance-constrained, and robust optimization [9], as shown in Fig. 2 Part B. Deterministic optimization enhances performance indicators under fixed conditions, boosting efficiency within a defined environment. In contrast, stochastic optimization considers uncertainty, optimizing expected performance across various scenarios. Chance-constrained optimization ensures LAENets meet predefined requirements with a probabilistic guarantee, managing error tolerance. Robust optimization, on the other hand, focuses on worst-case performance under uncertainty, ensuring LAENets operate safely in challenging conditions [9].
To optimize these robust indicators, discriminative AI methods typically are deterministic optimizations, which focus on learning decision boundaries. However, when uncertain influencing factors increase or fluctuate significantly, their performance tends to degrade rapidly. GenAI emphasizes modeling the underlying data distribution rather than decision boundaries, making it more effective for robust optimization.
@@ III GenAI-enabled Wireless Physical Layer for Robust LAENet
In this section, we describe the overview of GenAI. Then, we present some GenAI-enabled wireless physical layer communication methods and show how these methods improve the robust QoS metrics for LAENets.
@@ III-A Overview of GenAI
GenAI is a type of Artificial Intelligence (AI) that leverages unsupervised learning or self-supervised learning to generate new results that resemble the learned data, as depicted in Fig. 1 Part C. Typical generative AI models include Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and Diffusion Models (DMs), LLM [5]. GenAI excels in capturing and modeling specific data distributions, which allows for more accurate and detailed data analysis. For example, GenAI can detect subtle deviations that may indicate operational safety issues, enabling accurate threat detection. Additionally, it can generate synthetic data to simulate potential threats, such as spoofing attacks or signal interference in low-altitude systems, providing an adaptation mechanism [7]. These features can address robust issues encountered by LAENets, providing a solid foundation for leveraging GenAI models to enhance the robustness of LAENets in wireless physical layer communications.
III-B Channel Modeling
@@ III-D Beam Alignment
III-E Lessons Learned
Building on prior research, GenAI has shown significant potential in physical-layer wireless communications, especially in complex aerial scenarios. Its precise monitoring abilities enable aerial equipment to perform tasks accurately at high speeds, enhancing robust QoS metrics. GenAI employs a unique learning approach that maps latent distributions to key communication model parameters including path loss, delay, and angle measurements. This distributional perspective provides insights into correlated properties, enabling effective management of multiple parameters. However, current studies often rely on deterministic or stochastic optimizations, neglecting performance analysis under uncertainty conditions.
@@ IV-A Diffusion-based Wireless Optimization
GenAI models, particularly diffusion models combined with reinforcement learning, have demonstrated significant advantages in wireless network optimization [7]. These models inherently learn distributions, adapt to the effects of multi-scale fading in wireless channels, and exhibit a level of adaptability that traditional algorithms lack. For instance, the diffusion-based optimization framework GDM [13] employs a diffusion model-based actor to generate actions, as illustrated in Fig. 3 Part C. Through its multi-step denoising process, each step incrementally removes noise, enabling the actor to explore the entire state space and identify optimal actions. However, the denoising optimization process faces several challenges, particularly in uncertainty optimization. The stochastic nature of diffusion model generation can lead to increased variability in actions under the same state, affecting network convergence.
@@ IV-B MoE-transformer Actor Network
Inspired by this, we propose an MoE-transformer actor network structure to enhance the diffusion-based wireless optimization framework, as shown in Fig. 3 Part D. Our approach employs MoE layers to capture uncertainties across different states and denoising steps while integrating a transformer with a multi-head attention mechanism to stabilize generation. By replacing the action network in the diffusion reinforcement learning framework, the proposed actor network can utilize the advantages of the MoE structure and attention mechanism to achieve more stable and robust action generation. Therefore, the proposed method leverages the exploration and generative strengths of diffusion models to mitigate uncertainties in LAENet, enhancing overall system robustness.
@@ V-B Secure Beamformers
We consider a secure communication scenario where a terrestrial BS transmits a confidential signal along with artificial noise to communicate with legitimate UAVs and evade eavesdropping, as depicted in Fig. 3 Part A. The terrestrial BS employs a uniform planar array (UPA) with dimensions $N_{x}\times N_{y}$, whereas both legitimate UAVs and eavesdroppers (Eves) are equipped with uniform linear arrays (ULAs) comprising $N_{b}$ and $N_{e}$ antennas, respectively. Moreover, environmental factors, such as wind speed, will introduce estimation errors in the target UAV’s location. The covert nature of the eavesdropper further exacerbates uncertainty, while errors in channel estimation and angle-of-arrival estimation also influence beamforming accuracy, as shown in Fig. 3 Part B.
@@ V-C Optimization Design
To ensure reliable communication while mitigating eavesdropping risks, we accomplish secure beamforming by solving an optimization problem. Our objective is to maximize the achievable secrecy rate (ASR) by optimizing the beamformers for both the confidential signal and artificial noise, accounting for various sources of uncertainty. To restrict the eavesdropper’s ability, we impose a constraint that limits the capacity of the eavesdropping channel to remain below a specific threshold $C_{eve}$. We designed a reward function with a penalty item to ensure reinforcement learning algorithms can learn the optimal beamforming strategy under the constraints. Specifically, the reward function is the ASR minus the capacity of the eavesdropping channel over the threshold $C_{eve}$.
@@
  • Stochastic Optimization aims to maximize the mathematical expectation ASR for a given beamforming action under uncertainty. We incorporate the eavesdropping channel capacity as a penalty term to enforce constraints.
  • Chance-Constrained Optimization maximizes ASR while ensuring constraint satisfaction with a certain probability $P_{eve}$. When the constraint satisfaction probability is greater than $P_{eve}$, we only calculate the reward brought by ASR to maximize ASR from a probability perspective.
  • Robust Optimization focuses on guaranteeing worst-case performance under uncertainty. For a given beamforming action, we use the Monte Carlo method to approximate the worst-case with the lowest achievable reward. Then, we maximize this lower bound to enhance the robustness of the generated beamforming.
(a) Robust
(b) Normalized Inference Reward

### v2

@@ Abstract
Low-Altitude Economy Networks (LAENets) have emerged as significant enablers of social activities, offering low-altitude services such as the transportation of packages, groceries, and medical supplies. Owing to their control mechanisms and ever-changing operational factors, LAENets are inherently more complex and vulnerable to security threats than traditional terrestrial networks. As applications of LAENet continue to expand, the robustness of these systems becomes crucial. In this paper, we propose a generative artificial intelligence (GenAI) optimization framework that tackles robustness challenges in LAENets. We conduct a systematic analysis of robustness requirements for LAENets, complemented by a comprehensive review of robust Quality of Service (QoS) metrics from the wireless physical layer perspective. We then investigate existing GenAI-enabled approaches for robustness enhancement. This leads to our proposal of a novel diffusion-based optimization framework with a Mixture of Experts (MoE)-transformer actor network. In the robust beamforming case study, the proposed framework demonstrates its effectiveness by optimizing beamforming under uncertainties, achieving a more than 15% increase over four learning baselines in the worst-case achievable secrecy rate. These findings highlight the significant potential of GenAI in strengthening LAENet robustness.
@@ I Introduction
Recently, advancements in wireless communications and manufacturing technologies have propelled the development of drones and other aerial devices, greatly expanding their applications in low-altitude airspace. Enabled by these technologies, the Low-Altitude Economy (LAE) has emerged as a significant sector focusing on social opportunities within low-altitude airspace (defined as elevations up to 1,000 meters). The LAE aims to harness low-flying equipment for various social activities, including air traffic management, transportation, and surveillance, to organize the Low-Altitude Economy Networks (LAENet). These networks play an increasingly vital role by enabling essential services such as communication, computation, storage, and sensing [1]. According to the Civil Aviation Administration of China, the value of China’s low-altitude airspace economy exceeded CNY 500 billion in 2023 and is projected to reach CNY 3.5 trillion by 2035111https://www.swissre.com/institute/research/topics-and-risk-dialogues/china/insuring-low-altitude-airspace-economy-china.html. Moreover, the U.S. Federal Aviation Administration is set to finalize pilot training and certification rules for air taxis by 2024. A well-known research company, Mordor Intelligence, predicts that the urban air mobility market will reach approximately \$4.04 billion by 2029, demonstrating its significant social potential [^2] https://www.moomoo.com/news/post/44869062/faa-finalized-key-rules-usa-s-low-altitude-economy-leader.
@@
However, the unique low-altitude environment and the absence of unified management regulations have introduced several uncertainties to the design of LAENet. These uncertainties, including variable weather conditions, air traffic, user demand, and emergency situations, significantly heighten the requirement for LAENet devices to complete tasks in dynamic environments [1]. Therefore, robustness has been recognized as a Key Performance Indicator (KPI) that goes beyond efficiency (speed) and effectiveness (accuracy) for LAENet. Robustness refers to the ability to maintain the qualities of service (QoS) of LAENets despite the influence of uncertainties [2]. This recognition guides the evolution of LAENets, enabling innovation and commercial activity. For example, the tech giant Uber’s drone delivery utilizes advanced artificial intelligence algorithms, Global Positioning System (GPS) systems, and sensors to avoid obstacles under different weather conditions, ensuring maximum delivery safety [^3] https://karlobag.eu/en/technology/uber-launches-innovative-drone-food-delivery-service-in-urban-areas-using-high-technology-and-safety-measures-0tvcw. However, due to the complexity of the airspace environment, the high degrees of freedom, and the strong dynamics of aerial devices, building a robust wireless network platform for LAE applications is challenging.
@@
Enhancing the robustness of the wireless physical layer requires addressing uncertainty as a critical issue. However, both numerical algorithms and traditional learning methods face several challenges, including heavy dependence on training data, and susceptibility to unknown environments, which hinder their effectiveness in handling uncertainty. Generative artificial intelligence (GenAI) is an emerging technology that works to analyze data distribution. GenAI has demonstrated strong capabilities in enhancing wireless communications, including physical and network layers, through cross-dimensional data processing and generation [3]. Importantly, recent research highlights the potential of GenAI to capture the underlying distribution of optimal strategies, enhancing network optimizations [4]. Based on the achievements in wireless communications, there are several potential advantages of using GenAI to eliminate uncertainties and construct robust LAENets. GenAI can generate additional signal data, which is difficult to obtain in extreme scenarios, to handle diverse interference patterns. Additionally, GenAI can generate clean data, thereby removing the noise uncertainty in signals [3]. Moreover, GenAI enables optimization by modeling the distribution of strategies under uncertainties, allowing it to generate robust and improved solutions.
@@
Motivated by these analyses, this paper aims to investigate the utilization of GenAI to develop robust LAENets. The main contributions are summarized as follows:
  • We provide an overview of LAENets and their applications, offering a detailed analysis of the robustness within LAENet. Furthermore, we introduce various GenAI models within their applications and benefits in wireless communications toward LAENets.
  • From the perspective of the physical layer, we explore how GenAI enhances the robustness of wireless communications. Moreover, we present comprehensive robustness metrics and discuss the benefits of implementing GenAI models to support the construction of LAENets.
  • To systematically tackle robustness challenges, we incorporate robustness requirements into an optimization problem. We consider different levels of robustness through different optimization paradigms, including stochastic optimization, chance-constrained optimization, and robust optimization.
@@
II The Overview of Robust LAENet
@@
Figure 1: An overview of GenAI-enabled robust LAENet. Part A illustrates the architecture of LAENet from transportation, communication, and sensing aspects. Part B introduces three aspects of robustness in LAENets. Part C presents a comparison between GenAI and discriminative AI. Part D demonstrates important GAI-enabled applications, including channel modeling, channel estimation, and beam alignment.
@@ II-A Conceptions and Applications of LAENet
The LAE relies on the cooperation of edge equipment and low-altitude communication networks to facilitate social activities, which covers three primary aspects (shown in Fig. 1 Part A):
@@ II-B The Significance of Robust LAENet
  • Safety: Safety in LAENets involves ensuring the protection of equipment, operators, and service recipients during operations with low-altitude devices. Ensuring QoS of safety means that safety measures must consistently perform effectively, even when faced with challenges such as adverse weather, signal interference, or unexpected obstacles.
  • Security and Confidentiality: Security and confidentiality concerns LAENet’s ability to maintain secure and reliable low-altitude communications while ensuring QoS amidst various uncertainties. This involves resisting external attacks, protecting against data breaches, and preventing unauthorized access, all without compromising the performance required for different applications.
@@ II-C Robustness Considerations for LAENet
To effectively evaluate and enhance the robustness of LAENet, it is necessary to introduce robust QoS metrics in various applications. We first map key LAENet applications to suitable QoS indicators, transforming them into optimization problems under uncertainty, as illustrated in Fig. 2.
@@
Besides robust indicators, the effective adjustment and improvement of these indicators are typically realized through optimization methods. However, recognizing the varying nature of uncertainties inherent to LAENet operations, it is necessary to employ appropriate optimization frameworks to systematically address and manage these uncertainties. We introduce several optimization frameworks that offer increasing levels of robustness, including deterministic, stochastic, chance-constrained, and robust optimization [8], as shown in Fig. 2 Part B.
Deterministic optimization, which assumes a predictable environment, maximizes performance indicators under fixed conditions. It is suitable for scenarios where uncertainties are negligible. Conversely, stochastic optimization explicitly accounts for uncertainties by optimizing expected performance across a range of scenarios, making it effective in moderately uncertain environments. For scenarios with strict QoS requirements but limited error tolerance, chance-constrained optimization is employed. It ensures predefined requirements are met with a probabilistic guarantee, balancing performance and risk by controlling acceptable levels of constraint violation [8].
Robust optimization provides the highest level of reliability by optimizing worst-case performance, ensuring that LAENet systems remain operational under severe uncertainties or adversarial conditions. This method is particularly important in mission-critical applications where failure could result in significant consequences [8].
By aligning each robustness indicator with an appropriate optimization level, these optimization frameworks systematically enhance LAENet robustness, quantitatively improving critical QoS metrics under environmental uncertainties.
@@ III GenAI-enabled Wireless Physical Layer for Robust LAENet
In this section, we provide an overview of GenAI. Then, we present some GenAI-enabled wireless physical layer communication methods and show how these methods improve the robust QoS metrics for LAENets.
@@ III-A Overview of GenAI
GenAI is a type of Artificial Intelligence (AI) that leverages unsupervised learning or self-supervised learning to generate new results that resemble the learned data, as depicted in Fig. 1 Part C. Typical generative AI models include Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and diffusion models, LLM [3]. GenAI excels in capturing and modeling specific data distributions, which allows for more accurate and detailed data analysis. For example, GenAI can detect subtle deviations that may indicate operational safety issues, enabling accurate threat detection. Additionally, it can generate synthetic data to simulate potential threats, such as spoofing attacks or signal interference in low-altitude systems, providing an adaptation mechanism [6]. These features can address robust issues encountered by LAENets, providing a solid foundation for leveraging GenAI models to enhance the robustness of LAENets in wireless physical layer communications.
III-B Comparison of GenAI and discriminative AI
As shown in Fig. 1 Part C, GenAI can generate novel content by learning and mimicking the underlying data distributions. Distinct from traditional learning paradigms, such as discriminative AI, which are primarily designed to produce specific outcomes or decisions, GenAI excels in capturing and modeling specific data distributions, which allows for more accurate and detailed data analysis [7]. For instance, by learning the distributions of attacked and unattacked data, GenAI can analyze whether the input data conforms to these distributions, offering insights into potential anomalies and enhancing robustness, rather than merely making binary judgments about whether the input data has been attacked.
For optimization frameworks, discriminative AI methods traditionally employ deterministic optimization focusing primarily on decision boundaries. However, their performance degrades rapidly under high uncertainties [7]. In contrast, GenAI methods inherently model underlying data distributions, enabling a probabilistic representation of uncertainties that aligns naturally with stochastic, chance-constrained, and robust optimization paradigms. This probabilistic foundation allows GenAI to offer superior capabilities for robust optimization [4]. Additionally, GenAI addresses the limitations of traditional methods by generating additional synthetic data under extreme scenarios, removing noise uncertainties, and simulating realistic environmental conditions, thereby quantitatively enhancing robustness metrics such as BER, outage probability, and stability margins [3].
III-C Applications
III-C1 Channel Modeling
@@ III-C3 Beam Alignment
III-D Lessons Learned
Building on prior research, GenAI has shown significant potential in physical layer wireless communications, especially in complex aerial scenarios. Its precise monitoring abilities enable aerial equipment to perform tasks accurately at high speeds, enhancing robust QoS metrics. GenAI employs a unique learning approach that maps latent distributions to key communication model parameters including path loss, delay, and angle measurements. This distributional perspective provides insights into correlated properties, enabling effective management of multiple parameters. However, current studies often rely on deterministic or stochastic optimizations, neglecting performance analysis under uncertainty conditions.
@@ IV-A Diffusion-based Wireless Optimization
GenAI models, particularly diffusion models combined with reinforcement learning, have demonstrated significant advantages in wireless network optimization [7]. These models inherently learn distributions, adapt to the effects of multi-scale fading in wireless channels, and exhibit a high level of adaptability that traditional algorithms lack. For instance, the diffusion-based optimization framework GDM [4] employs a diffusion model-based actor to generate actions, as illustrated in Fig. 3 Part C. Through its multi-step denoising process, each step incrementally removes noise, enabling the actor to explore the entire state space and identify optimal actions. By contrast, other GenAI models, such as GAN-based or VAE actors, tend to either collapse to sub-optimal modes or struggle with the continuous, high-dimensional action spaces typical of beamforming and resource allocation problems [12]. Moreover, diffusion models provide exact log likelihoods. Because GANs have no explicit density and VAE actors introduce approximation error, diffusion-based policies deliver more reliable uncertainty estimates and can adapt quickly to sudden channel variations [12].
However, the denoising optimization process faces several challenges, particularly in uncertainty optimization. The stochastic nature of diffusion model generation can lead to increased variability in actions under the same state, affecting network convergence. Therefore, our ongoing work augments the actor network within the diffusion-based optimization framework to achieve a balance between exploration and stability, further justifying diffusion models as a principled and practically effective choice for robust wireless optimization.
@@ IV-B MoE-transformer Actor Network
In LAENets, the rapid variations in channel conditions, traffic loads, and adversarial threats significantly increase system uncertainties. Multi-head attention mechanisms are particularly suitable for modeling fine-grained state dependencies and representing these uncertainties effectively on the fly. An MoE layer retains this attention expressiveness while sparsely activating only the most relevant experts, slashing computing and energy demands to meet the tight resource availability of LAE edge nodes [7]. By integrating the two into an MoE-transformer network, the model effectively captures dynamic uncertainties while scaling computation adaptively, achieving the accuracy-efficiency trade-off essential for LAENets.
Inspired by this, we propose an MoE-transformer actor network structure to enhance the diffusion-based wireless optimization framework, as shown in Fig. 3 Part D. Our approach integrates a transformer with multi-head attention to capture fine-grained dependencies and dynamic uncertainties across states and denoising steps, while MoE layers selectively activate the most relevant experts to stabilize generation and reduce computation. The proposed actor network is specifically designed within the diffusion reinforcement learning framework to utilize the strengths of the MoE structure and attention mechanism, enabling more stable and robust action generation. Therefore, the proposed method leverages the exploration and generative strengths of diffusion models to mitigate uncertainties in LAENets, enhancing overall system robustness.
@@ V-B Secure Beamformers
We consider a secure communication scenario where a terrestrial BS transmits a confidential signal along with artificial noise to communicate with legitimate UAVs and evade eavesdropping, as depicted in Fig. 3 Part A. The terrestrial BS employs a uniform planar array (UPA) with dimensions $N_{x}\times N_{y}$, whereas both legitimate UAVs and eavesdroppers (Eves) are equipped with uniform linear arrays (ULAs) comprising $N_{b}$ and $N_{e}$ antennas, respectively. Moreover, environmental factors, such as wind speed, will introduce estimation errors in the target UAV’s location. The covert nature of the eavesdropper further exacerbates uncertainties, while errors in channel estimation and angle-of-arrival estimation also influence beamforming accuracy. As illustrated in Fig. 3 Part B, these uncertainties are incorporated in the optimization input. In simulations, we emulate varying uncertainty levels by injecting controlled noise.
@@ V-C Optimization Design
To ensure reliable communication while mitigating eavesdropping risks, we accomplish secure beamforming by solving an optimization problem. Our objective is to maximize the achievable secrecy rate (ASR) by optimizing the beamformers for both the confidential signal and artificial noise, accounting for various sources of uncertainties. To restrict the eavesdropper’s ability, we impose a constraint that limits the capacity of the eavesdropping channel to remain below a specific threshold $C_{eve}$. We designed a reward function with a penalty item to ensure reinforcement learning algorithms can learn the optimal beamforming strategy under the constraints. Specifically, the reward function is the ASR minus the capacity of the eavesdropping channel over the threshold $C_{eve}$.
@@
  • Stochastic Optimization aims to maximize the mathematical expectation ASR for a given beamforming action under uncertainties. We incorporate the eavesdropping channel capacity as a penalty term to enforce constraints.
  • Chance-Constrained Optimization maximizes ASR while ensuring constraint satisfaction with a certain probability $P_{eve}$. Since direct probability computation is intractable, we employ a Monte Carlo method to approximate constraint satisfaction probability for all reinforcement learning methods. When the constraint satisfaction probability is greater than $P_{eve}$, we only calculate the reward brought by ASR to maximize ASR from a probability perspective.
  • Robust Optimization focuses on guaranteeing worst-case performance under uncertainties. For a given beamforming action, we use the Monte Carlo method to approximate the worst-case with the lowest achievable reward. Then, we maximize this lower bound to enhance the robustness of the generated beamforming.
(a) Stochastic
(b) Chance-constrained
(c) Robust
(d) Inference Reward